5. Home for less Corporation's bonds have a 10-year maturity, a 6.25% semiannual coupon, and a par value of $1,000. The going interest rate (rd) is 4.75%, based on semiannual compounding. What is the bond’s price?
1
Expert's answer
2020-06-17T10:28:50-0400

"P=\frac{1000*0.0625*(1/1.0475^{10}-1)}{1/1.0475-1}+\frac{1000}{1.0475^{10}}=1140.45"